OpenStudy (anonymous):

Given: F(x) = x + 2 and G(x) = 3x + 5 (F - G) (x) =

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

(F - G) (x) = x+2-(3x+5) = x+2-3x-5 = ...

OpenStudy (anonymous):

-3

OpenStudy (raden):

combine the similar terms, = (x-3x) + (2-5) = -2x - 3

OpenStudy (anonymous):

Given: f(x) = x + 2 [f(x)] 2 =